# Bridging Synthetic and Real Worlds for Pre-training Scene Text Detectors

### v1

@@ 3.3 Discussion on Character Region Awareness
As illustrated in Fig. 3, existing LSD only contains English texts with tightly spaced characters, while URD may encompass text properties from various languages (such as Japanese, Korean, and Chinese). Consequently, even though the model predictions effectively cover text regions, we encounter challenges when parsing these predictions into word-level bounding boxes or polygons.
Indeed, when using English as the source domain and adapting it to various language domains, it appears that no matter how sophisticated the models and loss functions are, they are unable to accurately and effectively parse the complex text layout and semantic context. This challenge primarily arises from variations in the spacing and arrangement of characters, words, and sentences across languages. This observation leads us to revisit the learning objective of scene text detectors during pre-training.
We emphasize an alternative perspective – utilizing characters as the fundamental learning unit for text detection. The strategy presents several advantages: 1) Like word-level bounding boxes (WBB), character-level bounding box (CBB) annotations can be readily acquired from the existing synthetic dataset (SynthText) at no additional cost; 2) Instead of parsing complex layouts and spacing disparities across languages during pre-training, decoding them into individual characters is straightforward and attainable; 3) The pre-trained model with character-level region awareness easily integrates these learned character features during fine-tuning to adapt text attribute rules provided by different language text datasets.
To the end, we select CBB annotations as $\mathcal{B}_{l}$. Following the pipeline outlined in Fig. 4, the mixed image $\mathbf{X}_{g}$ is created by cropping the synthetic image $\mathbf{X}_{l}$ using CBB annotations $\mathcal{B}_{l}$ and pasting its glyphs onto the unlabeled image $\mathbf{X}_{u}$. The corresponding mixed pseudo labels $\mathbf{Y}_{g}$ are then obtained by combining the pixel-wise binary map $\mathbf{Y}_{l}$ which is converted from $\mathcal{B}_{l}$, with the estimated character-level pseudos $\mathbf{Y}_{u}$ generated by the teacher branch.
@@ 4.1 Dataset
Unlabeled Real Data (URD). Our collected URD is sourced from a wide array of natural scenes, encompassing a diverse range of fonts, languages, text shapes, and sizes. Specifically, URD inlcudes 430K images from LSVT [32], 9K from MLT17 [23], 19657 from MLT19 [24], 229 from ICDAR2013 [14], and 500 from USTB-SV1K [49].
Benckmarks. Our method is designed to detect text with arbitrary shapes, including oriented and curved texts from the following publicly available benchmarks:
Total-Text (TT) [5] consists of 1255 training images and 300 test images. It contains many curved text instances with polygon-shaped bounding box annotations.
CTW 1500 (CTW) [19] mainly focuses on curved texts and consists of 1000 training images and 500 test images with polygon-shaped bounding box annotations.
ICDAR2015 (IC15) [15] contains 1000 training images and 500 testing images with word-level rectangular bounding box annotations.
MSRA-TD500 (TD) [46] includes 300 training images and 200 test images with line-level rectangular bounding box annotations. HUST-TR400 dataset [47] is typically added to train the model.
@@ 4.2 Implementation Details
Referring to recent works [34, 31, 44, 50], we combine representative text detection methods, including DBNet [17], PANet [38], PSENet [37], and FCENet [57], to conduct text detection experiments. All experiments are implemented on a server with 2 NVIDIA A800 GPUs, using PyTorch.
Pre-Training. The pre-training experiments are conducted on labeled synthetic data (LSD) and unlabeled real data (URD), utilizing the open-source codes provided by these text detection methods. For synthetic images, we adopt consistent augmentation settings as prescribed by these methods for a fair comparison. For unlabeled real images, we employ a standard weak-to-strong augmentation technique [56] to enrich diversity. Our method is trained using the AdamW optimizer [22], along with a cosine learning rate scheduler [6]. This includes a warm-up phrase for 1 epoch in a total of 10 epochs. The base learning rate is set to $2560.003\times\rm{batchsize}/$ 256. At each iteration, we equally sample LSD and URD with a total batch size of 24.
Fine-Tuning. For a fair comparison, we employ the same fine-tuning configurations as these detection methods and then evaluate them on these publicly available benchmarks: TT, CTW, IC15 and TD, following the standard metric: Precision (P), Recall (R), and F-measure (F). The teacher model, chosen from the final pre-training epoch, serves as the corresponding weights for the fine-tuning stage.
@@ 4.3 Experiment Results
Table 1: Comparison with SOTA pre-training methods for text detectors on TT, CTW, IC15, and TD. F-measure (%) is reported. ${}^{\divideontimes}$ denotes the use of extra labeled real data. These symbols $\dagger$, ${}^{\ast}$, $\ddagger$,  and ${}^{\star}$ indicate the results from  [51],  [31],  [44] and  [36], respectively, within the scope of this paper. The best results are shown in bold font. The second-best results are underlined.
@@
${}^{\divideontimes}$  [48] CVPR’23 87.6 - 89.8 -
${}^{\divideontimes}$  [12] ICCV’23 90.0 - 91.0 89.5
SIR [27] MM’23 88.2 85.5 87.8 89.6
@@
${}^{\star}$ [36] MM’23 85.7 - - 85.6
@@
DBNet+FreeReal(ours) - 88.9 87.9 90.0 90.1
@@
PANet+FreeReal (ours) - 87.4 88.4 86.0 89.9
4 ${}^{\ast}$ [37] CVPR’19 83.9 81.6 81.3 -
${}^{\ast}$ [34] CVPR’21 84.3 82.9 83.7 -
${}^{\ast}$ [31] CVPR’22 86.1 83.3 84.3 -
[…]
PSENet+FreeReal (ours) - 86.8 86.1 85.2 86.0
[…]
FCENet+FreeReal (ours) - 87.9 87.2 87.7 89.0
To ensure a comprehensive comparison, we collect results from public evaluation benchmarks reported by existing pre-training methods, and organize them in Table 1. The sources of these results are clearly recorded in the table’s caption using different symbols for easy reference. Specifically, these methods can be classified into two pre-training paradigms. In Paradigm A, models (i.e., +ST, +TCM, +FastTCM) are initially pre-trained on the synthetic dataset and then fine-tuned on real datasets. In Paradigm B, models (i.e., +STKM, +VLPT, +oCLIP, +MTM) are first pre-trained by various pretext tasks, and the well-learned backbone weights are transferred into scene text detectors when fine-tuning on real datasets.
1) Results on DBNet. As indicated in the second group of rows in Table 1, all pre-training methods both report their cooperative performance with DBNet. When compared to methods using paradigm A, differing from DBNet+ST, we exclude the use of a deformable convolutional network (DCN) in our model. Despite these differences, our method still achieves impressive results, yielding improvements of 4.2%, 4.5%, 4.6%, and 5.2% on TT, CTW, IC15, and TD datasets, respectively. Compared to DBNet+TCM and DBNet+FastTCM methods, our method achieves SOTA detection performance on four datasets, particularly on curved text datasets (TT and CTW), with gains of 2.8% and 2.7%, respectively. In comparison with methods following Paradigm B, our proposed method consistently outperforms previous SOTA works by substantial margins, achieving performance improvements of 2.6%, 5.9%, 3.5% and 1.6% on TT, CTW, IC15, and TD datasets, respectively. In general, when compared to all methods, our proposed method exhibits effectiveness and robustness.
2) Results on PANet. As shown in the third group of rows in Table 1, our method outperforms PANet+ST on four benchmarks, achieving a gain of 3.85% in terms of average F-measure, while using the same parameters and labeled data for pre-training and fine-tuning. Furthermore, our method exhibits its prominent superiority, particularly on the CTW and TD datasets, surpassing the PANet+FastTCM method by a substantial margin of 3.9% and 4.5%, respectively.
3) Results on PSENet. Based on the observation from the fourth group of rows from Table 1, only methods using paradigm B report the performance in collaboration with PSENet. In comparison to these methods, our proposed method brings 0.7%, 2.8%, and 0.9% performance gains on TT, CTW, and IC15 datasets, respectively.
4) Results on FCENet. In the last group of rows from Table 1, we conduct a comparative experiment with pre-training methods in collaboration with FCENet, a text detector that operates by regressing Fourier signature vectors derived from ground-truth point sequences. It can be observed that our method further refreshes the best text detection results and improves 2.1%, 1.2%, 0.4% and 1.9% on TT, CTW, IC15, and TD datasets, respectively. These experimental results demonstrate the robustness of our method, even when cooperated with regression-based methods.
Table 2: Comparison with existing pre-training paradigms when fine-tuning on the TotalText dataset. These symbols $\dagger$, $\ddagger$,  and ${}^{\star}$ indicate the results from  [51],  [44] and  [36], respectively. $\Delta$F is the improvement relative to the baseline.
Paradigm Method Data F-measure $\Delta$ F(%)
@@
[…]
${}^{\star}$ [36] 800K 41K 85.7 +7.5%
C Ours (10%) 80K 46K 87.5 +9.3%
Ours (20%) 160K 92K 88.1 +9.9%
Ours (50%) 400K 230K 88.3 +10.1%
Ours (100%) 800K 460K 88.9 +10.7%
5) Paradigm Comparison. In this section, we aim to investigate which kinds of pre-training paradigms for text detectors are important and meaningful for future studies. To this end, we conduct a comparative analysis with two mainstream pre-training paradigms mentioned above, as summarized in Table 2. When compared to these well-curated designs at the cost of introducing additional auxiliary pretext tasks or more training modules, our proposed paradigm achieves a significant improvement of 1.2% (87.5 vs. 86.3), despite utilizing 10% of the labeled synthetic data and 10% of the unlabeled real data. This outcome underscores our ability to harness the potential of unlabeled real images. Moreover, when we gradually increase the proportion of the total training data, the text detection performance continues to improve. Specifically, our method achieves gains of 1.8%, 2.0%, and 2.6% at data fractions of 20%, 50%, and 100%, respectively, compared to the recent SOTA method. This outcome fully demonstrates the superiority of our pre-training paradigm, which integrates the complementary strengths of both LSD and URD for pre-training.
6) Results on Unseen Language Text. Our pre-trained model is capable of learning the basic character feature representations of languages and then integrating them during fine-tuning. To show its robustness, we choose to evaluate on unseen language text dataset VinText [25]. FreeReal achieves an improvement of 0.7%, as shown in Table 3.
Table 3: Comparison with supervised SOTA work on unseen language text. “R50” and “DeT” denote a ResNet50 and Deformable Transformer structure. ${}^{\divideontimes}$ denotes the use of extra labeled real data.
@@
Table 4: Evaluate the effectiveness of the proposed modules.  $\Delta$F is the improvement of F-measure relative to baseline. “Stu-Tea”, “CRA”, “GlyphMix” denote the standard student-to-teacher framework, using character-level bounding boxes (CBB), and the glyph-based mixing mechanism, respectively.
Stu-Tea
CRA
GlyphMix $\Delta$ F(%)
P
R
F
$\times$
$\times$
$\times$ -
87.1
82.5
84.7
$\times$
$\times$ +1.31%
88.80
83.39
86.01
$\times$ +1.73%
87.76
85.15
86.43
$\times$
+3.12%
89.10
86.58
87.82
+3.40%
88.81
87.40
88.10
5 Ablations and analysis
For efficiency, all ablation experiments are implemented in collaboration with DBNet method [17]. This involved pre-training with 20% of LSD and 20% of URD, followed by evaluation on the curved TotalText (TT) dataset.
Effectiveness of Network Components. To comprehensively evaluate the effectiveness of our proposed method, we divide it into three components and evaluate each one individually, as shown in Table 4. First, we enhance the baseline DBNet model by incorporating a standard student-teacher framework, yielding a performance gain of 1.31%. Second, within the student-teacher framework, we evaluate the individual contributions of the CRA and GlyphMix components, achieving 1.73% and 3.12% performance improvements, respectively. Finally, upon combining all the plug-ins, the overall performance reaches 88.1% F-measure, showing a substantial gain of 3.40% compared to the baseline. These experiments consistently demonstrate our proposed method effectively bridges both the synth-to-real and language-to-language domain gaps when leveraging the intrinsic qualities of any unlabeled real data.
@@ 5 Ablations and analysis
Table 5: Evaluate the effectiveness of GlyphMix components. “GIM” and “TIM” denote the Graffiti-like Inter-domain Mixing and Text-balanced Intra-domain Mixing, respectively.
GIM
TIM
$\Delta$ F(%) P
R
F
$\times$
$\times$
+1.73% 87.76
85.15
86.43
$\times$
+2.95% 88.64
86.68
87.65
$\times$
+3.01% 88.29
87.13
87.71
Effectiveness of GlyphMix Components. In Table 5, we show the performance gains brought by the two components of GlyphMix, TIM and GIM, respectively.
@@
method
DCA P R F $\Delta$ F(%)
Stu-Tea
- 88.80 83.39 86.01 -
Mixup[55]
69.5 89.17 84.12 86.57 +0.56%
CutMix[52]
72.8 87.99 85.64 86.80 +0.79%
ClassMix[26]
78.9 90.65 83.61 86.99 +0.98%
FreeReal (ours)
95.7 88.81 87.40 88.10 +2.09%
Compared with Bridging-domain Methods. In this section, we begin with a domain classification experiment, calculating the confidence that the mixed images generated by domain bridging methods belong to the real domain. To accomplish this, we employ a network architecture, consisting of resnet18 as a feature extractor, followed by a binary classification head, for simplicity. The primary objective is to distinguish between synthetic and real images by assigning them to different classes. Subsequently, we evaluate the performance of the well-established network in different domain bridging methods, as summarized in Table 6. Remarkably, when compared to these methods, GlyphMix achieves an outstanding performance score of 95.7%, surpassing the second-best method by a substantial margin of 16.8%. This result demonstrates that our method is able to generate arbitrarily realistic images without domain drift.
Compared with Semi-supervised Methods. In this section, we compared the performance with recent SSL methods. These methods tend to adopt increasingly complex designs, involving advanced contrastive learning techniques, additional network components, multiple ensemble models, and complex training procedures. However, some designs, such as Gentle Teaching Assistant, may not be conducive to text feature learning because they do not fully consider the noticeable domain gap in text images. In contrast, we aim to eliminate domain gaps between LSD and URD. Despite adopting the standard student-teacher framework, our method still exceeds the previous SOTA by 0.88% F-measure value in Table 7.
@@
Effectiveness of $\mathbf{M}_{G}$. In this section, we evaluate the accuracy of character segmentation results $\mathbf{M}_{G}$ using TotalText [5] and TextSeg [43] datasets, both of which provide pixel-level annotations. Specifically, compared to conducting the cluster task on the whole image, we crop it into several text instances and then perform clustering to form $\mathbf{M}_{G}$. While the former achieves 32.35% and 45.54% mIoU, we report the mIoU of $\mathbf{M}_{G}$, i.e., 80.43% for TotalText and 84.14% for TextSeg, respectively. Notably, the underlying morphological representations of glyphs exhibit relative invariance to slight structural changes, such as thicker or thinner. This reduces the dependence on pixel-level high-precision segmentation with expensive costs. Furthermore, although the clustering process introduces some noise, leading to inaccuracies in $\mathbf{M}_{G}$, it can still be viewed as a way of adding data noise in the image-mixing process, enhancing the robustness of text detection.

### v3

@@ 3.2 Discussion on Character Region Awareness
Existing LSD only contains English texts with tightly spaced characters, while URD may encompass text properties from various languages (such as Japanese, Korean, Chinese, etc). Consequently, even though the model predictions effectively cover text regions, we encounter challenges in parsing these predictions into word-level bounding boxes or polygons.
Indeed, when using a single language (e.g., English) as the source domain and adapting it to various language domains, it appears that no matter how sophisticated the models and loss functions are, they struggle to accurately and effectively parse the complex text layout and semantic context. This challenge primarily arises from variations in the spacing and arrangement of characters, words, and sentences across languages. This observation leads us to revisit the learning objective of scene text detectors during pre-training.
We emphasize an alternative perspective utilizing characters as the fundamental learning unit for text detection. The strategy presents several advantages: 1) Like word-level bounding boxes (WBB), character-level bounding box (CBB) annotations can be readily acquired from the existing synthetic dataset (SynthText) at no additional cost; 2) Instead of parsing complex layouts and spacing disparities across languages during pre-training, decoding them into individual characters is straightforward and attainable; 3) The pre-trained model with character-level region awareness easily integrates its learned character features to adapt text attribute rules provided by different language text datasets during fine-tuning. To this end, we select CBB annotations as $\mathcal{B}_{l}$.
@@ 4.1 Dataset
Unlabeled Real Data (URD). URD is collected from a wide array of natural scenes, encompassing a diverse range of fonts, languages, and text shapes. Specifically, URD includes 430K images from LSVT [39], 9K from MLT17 [28], 19657 from MLT19 [27], 229 from ICDAR2013 [17], and 500 from USTB-SV1K [57].
Benchmarks. Four publicly available benchmarks are used to evaluate the per-
formance, including 1) Total-Text (TT) [6], 2) CTW 1500 (CTW) [23], 3) ICDAR2015 (IC15) [16], and 4) MSRA-TD500 (TD) [55].
@@ 4.2 Implementation Details
Referring to recent works [41, 37, 52], we combine representative text detection methods [20, 46, 44, 66] to conduct experiments. URD is used in Tabs. 1-3 and the same data is used in Tabs. 4-7. All experiments are implemented on a server with 2 NVIDIA A800 GPUs, using PyTorch. 1) Pre-Training. The pre-training experiments are conducted on LSD and URD, utilizing the open-source codes provided by these text detection methods. For synthetic images, we adopt consistent augmentation settings as prescribed by these methods for a fair comparison. For unlabeled real images, we employ a standard weak-to-strong augmentation technique [64] to enrich diversity. Our method is trained using the AdamW optimizer [26], along with a cosine learning rate scheduler [7]. This includes a warm-up phrase for 1 epoch in a total of 10 epochs. The base learning rate is set to $2560.003\times\rm{batchsize}/$ 256. At each iteration, we equally sample LSD and URD with a total batch size of 24. 2) Fine-Tuning. For a fair comparison, we employ the same fine-tuning configurations as these detection methods and then evaluate them on these publicly available benchmarks: TT, CTW, IC15 and TD, following the standard metric: Precision (P), Recall (R), and F-measure (F).
@@ 4.3 Experiment Results
To ensure a comprehensive comparison, we collect results from public evaluation benchmarks reported by existing pre-training methods, and organize them in Table 1. The sources of these results are clearly recorded in the table’s caption using different symbols for easy reference. Specifically, these methods can be classified into three pre-training paradigms. In Paradigm A, models (i.e., +ST, +TCM) are initially pre-trained on LSD and then fine-tuned on real datasets directly. In Paradigm B, models (i.e., +STKM, +VLPT, +oCLIP) are first pre-trained on LSD by various pretext tasks, and the well-learned backbone weights are transferred into scene text detectors when fine-tuning on real datasets. In Paradigm C, models (i.e., +MTM) are pre-trained on URD by pretext tasks and on LSD by supervised learning, and then fine-tuned on real datasets.
1) Results on DBNet. As indicated in the second group of rows in Table 1, all pre-training methods both report their cooperative performance with DBNet. When compared to methods using paradigm A, differing from DBNet+ST, we exclude the use of a deformable convolutional network (DCN) in our model. Despite these differences, our method still achieves impressive results, yielding improvements of 4.2%, 4.5%, 4.6%, and 5.2% on TT, CTW, IC15, and TD datasets, respectively. Compared to DBNet+TCM methods, our method achieves SOTA detection performance on four datasets, particularly on curved text datasets (TT and CTW), with gains of 3.0% and 2.8%, respectively. In comparison with methods following Paradigm B, our method consistently outperforms previous SOTA works by substantial margins, achieving performance improvements of 2.6%, 5.9%, 3.5% and 1.6% on TT, CTW, IC15, and TD datasets, respectively. In comparison with DBNet+MTM following Paradigm C, our method gets 3.2% and 4.5% gains on TT and TD datasets, respectively.
Table 1: Comparison with SOTA pre-training methods for text detectors on TT, CTW, IC15, and TD. F-measure (%) is reported. denotes the use of extra labeled real data. For easy reference, we mark these symbols $\dagger$, , $\ddagger$,  and separately to indicate the results from [59], [37], [52], and [43], within the scope of this paper. The best results are shown in bold font. The second-best results are underlined.
@@
DeepSOLO  [56] CVPR’23 87.6 - 89.8 -
ESTextSpotter  [14] ICCV’23 90.0 - 91.0 89.5
SIR [31] ACM MM’23 88.2 85.5 87.8 89.6
@@
DBNet+MTM [43] ACM MM’23 85.7 - - 85.6
DBNet+FreeReal (ours) - $\mathbf{88.9}_{(+2.6)}$ $\mathbf{87.9}_{(+2.8)}$ $\textbf{90.0}_{(+0.6)}$ $\mathbf{90.1}_{(+1.3)}$
@@
[…]
PANet+FreeReal (ours) - $\mathbf{87.4}_{(+2.4)}$ $\mathbf{88.4}_{(+4.1)}$ $\mathbf{86.0}_{(+1.4)}$ $\mathbf{89.9}_{(+4.6)}$
4 PSENet+ST [44] CVPR’19 83.9 81.6 81.3 -
PSENet+STKM [41] CVPR’21 84.3 82.9 83.7 -
PSENet+VLPT [37] CVPR’22 86.1 83.3 84.3 -
[…]
PSENet+FreeReal (ours) - $\mathbf{86.8}_{(+0.7)}$ $\mathbf{86.1}_{(+2.8)}$ $\mathbf{85.2}_{(+0.9)}$ 86.0
@@
FCENet+FreeReal (ours) - $\mathbf{87.9}_{(+2.1)}$ $\mathbf{87.2}_{(+1.3)}$ $\mathbf{87.7}_{(+0.6)}$ $\mathbf{89.0}_{(+2.1)}$
2) Results on PANet. As shown in the third group of rows in Table 1, our method outperforms PANet+ST on four benchmarks, achieving a gain of 3.85% in terms of average F-measure, while using the same parameters and labeled data for pre-training and fine-tuning. Furthermore, our method exhibits its prominent superiority, particularly on the CTW and TD datasets, surpassing the PANet+TCM method by a substantial margin of 4.1% and 4.6%, respectively.
3) Results on PSENet. FreeReal brings 0.7%, 2.8%, and 0.9% gains on TT, CTW, and IC15 datasets, respectively.
4) Results on FCENet. FreeReal further refreshes the best text detection results and improves 2.1%, 1.3%, 0.6% and 2.1% on TT, CTW, IC15, and TD datasets, respectively. These experimental results demonstrate the robustness of FreeReal, even when cooperated with regression-based methods.
Table 2: Comparison with existing pre-training paradigms when fine-tuning on the TotalText dataset. These symbols $\dagger$, $\ddagger$,  and indicate the results from [59], [52] and [43], respectively. $\Delta$F is the improvement relative to the baseline.
Paradigm Method Venue Data F-measure $\Delta$ F(%)
@@
C DBNet+MTM [43] ACM MM’23 800K 41K 85.7 +7.5%
D Ours (9%) - 71K 41K 87.4 +9.2%
Ours (10%) - 80K 46K 87.5 +9.3%
Ours (20%) - 160K 92K 88.1 +9.9%
Ours (50%) - 400K 230K 88.3 +10.1%
Ours (100%) - 800K 460K 88.9 +10.7%
5) Paradigm Comparison. In this section, we aim to investigate which kinds of pre-training paradigms for text detectors are important and meaningful for future studies. To this end, we conduct a comparative analysis with three mainstream pre-training paradigms mentioned above, as summarized in Table 2. When compared to these well-curated designs at the cost of introducing additional auxiliary pretext tasks or more training modules, our proposed paradigm achieves a significant improvement of 1.2% (87.5 vs. 86.3), despite utilizing 10% of the labeled synthetic data and 10% of the unlabeled real data. This outcome underscores our ability to harness the potential of unlabeled real images. Moreover, when we gradually increase the proportion of the total training data, the text detection performance continues to improve. Specifically, our method achieves gains of 1.8%, 2.0%, and 2.6% at data fractions of 20%, 50%, and 100%, respectively, compared to the recent SOTA method. This outcome fully demonstrates the superiority of our pre-training paradigm, which jointly leverage LSD and URD by bridging their gap for enhancing the text detector pre-training.
7) Results on Unseen Language Text. Our pre-trained model is capable of learning the basic character feature representations of languages and then combining them during fine-tuning. To show its robustness, we choose to evaluate on unseen language text dataset VinText [29]. FreeReal achieves an improvement of 0.7%, as shown in Table 3.
5 Ablations and analysis
Unless otherwise stated, all ablation experiments are implemented in collaboration with DBNet method [20] using 20% of LSD and 20% of URD and evaluated
on the curved TotalText (TT) dataset for efficiency.
Effectiveness of Network Components. To comprehensively evaluate the effectiveness of our proposed method, we divide it into three components and evaluate each one individually, as shown in Table 4. “Stu-Tea”, “CRA”, “GlyphMix” denote the standard student-to-teacher framework, Character Region Awareness introduced in Sec.3.2 to bridge the linguistic domain gap, and the glyph-based mixing mechanism, respectively. First, we enhance the baseline DBNet model by incorporating a standard student-teacher framework, yielding a performance gain of 1.31%. Second, within the student-teacher framework, we evaluate the individual contributions of the CRA and GlyphMix components, achieving 1.73% and 3.12% performance improvements, respectively. Finally, upon combining all plug-ins, the overall performance reaches 88.1%, showing a substantial gain of 3.40% compared to the baseline. These experiments consistently demonstrate that FreeReal effectively bridges both the synth-to-real and language-to-language domain gaps when leveraging the intrinsic qualities of unlabeled real data.
Table 3: Comparison with supervised SOTA work on unseen language text. “R50" and “DeT" denote a ResNet50 and Deformable Transformer structure, respectively. denotes the use of extra labeled real data.
@@ 5 Ablations and analysis
Table 4: Evaluate the effectiveness of the proposed modules. $\Delta$F is the improvement of F-measure relative to baseline.
Stu-Tea CRA GlyphMix $\Delta$ F(%) P R F
$\times$ $\times$ $\times$ - 87.1 82.5 84.7
$\times$ $\times$ +1.31% 88.80 83.39 86.01
$\times$ +1.73% 87.76 85.15 86.43
$\times$ +3.12% 89.10 86.58 87.82
+3.40% 88.81 87.40 88.10
Table 5: Evaluate the effectiveness of GlyphMix components. “GIM" and “TIM" denote the Graffiti-like Inter-domain Mixing and Text-balanced Intra-domain Mixing.
GIM TIM $\Delta$ F(%) P R F
$\times$ $\times$ - 87.76 85.15 86.43
$\times$ +1.22% 88.64 86.68 87.65
$\times$ +1.28% 88.29 87.13 87.71
@@
Comparison with pre-training methods using URD. Bridging the synth-to-real gap to simultaneously leverage LSD and URD for pre-training is our major contribution. This capability, not present in other PSTD methods, enables seamless incorporation of the huge number of URD to enhance pre-training. Thus to compare to other PSTD methods with complex pretext task designs, we use different ratios of URD to explore the effectiveness of our pre-training paradigm. Even when using 10% LSD and 10% URD in Table 2, our method still achieves SOTA with a 1.2% gain, which exhibits our method can serve as a novel baseline for PSTD. Besides, we add the comparison using additional URD to train other PSTD methods. In this case, these mentioned methods can be viewed the same as MTM [43]. As a result, FreeReal outperforms MTM by 1.7% in Table 2.
Effectiveness of GlyphMix Components. In Table 5, GlyphMix comprises two modules, inter-domain GIM and intra-domain TIM, contributing gains of 1.28% and 1.22% to the final performance, respectively. Besides, when compared to the data engine tool [11], GIM achieves intra-domain mixing in a cost-effective way, which yields real-world images online via K-means, a more flexible, faster and direct process than the multi-stage, pre-generated method. Specifically, GIM attains a 10.7x speedup (2.6h vs 27.7h) in generating 92K images, which is more suitable for real-time online training tasks. Additionally, GIM can avoid improper text blends on URD by leveraging the learned knowledge (boxes/maps) from our pre-training framework online. We also conduct the ablations and get 0.5% (85.3% vs 84.8%) performance gains, when using 5% training data.
@@
method DCA P R F $\Delta$ F(%)
Stu-Tea - 88.80 83.39 86.01 -
Mixup[63] 69.5 89.17 84.12 86.57 +0.56%
CutMix[60] 72.8 87.99 85.64 86.80 +0.79%
ClassMix[30] 78.9 90.65 83.61 86.99 +0.98%
FreeReal (ours) 95.7 88.81 87.40 88.10 +2.09%
@@
Comparison with Bridging-domain Methods. The goal of generating mixed images through domain-bridging methods is to bring synthetic images closer to real images. Consequently, the adaptation capability can be viewed as the likelihood of mixed images belonging to the real domain. A classification task is employed to assess the capability, whose the evaluation protocol is defined: Given a well-learned domain classifier $\phi$ (trained by n synthetic images $x_{l}\in X_{L}$ and n real images $x_{u}\in X_{U}$) and domain bridging method $\xi$, the classifier predicts the domain label (0 for synthetic and 1 for real) of mixed images and evaluates the accuracy as $\frac{\sum_{x_{l}\in X_{L},x_{u}\in X_{U}}\phi(\xi(x_{l},x_{u}))}{n^{2}}$. To accomplish this, we employ a network architecture, consisting of resnet18 as a feature extractor, followed by a binary classification head, for simplicity. Subsequently, the performance of different domain bridging methods is summarized in Table 6. Remarkably, when compared to these methods, GlyphMix achieves an outstanding performance score of 95.7%, surpassing the second-best method by a substantial margin of 16.8%. This result demonstrates that our method is able to generate arbitrarily realistic images without domain drift.
Comparison with Semi-supervised Methods. In this section, we compared the performance with recent SSL methods. These methods tend to adopt increasingly complex designs, involving advanced contrastive learning techniques, additional network components, multiple ensemble models, and complex training procedures. In contrast, we aim to eliminate domain gaps between LSD and URD. Despite adopting the standard student-teacher framework, our method still exceeds the previous SOTA by 0.88% in Table 7.
Effectiveness of $\mathbf{M}_{G}$. We evaluate the accuracy of character segmentation results $\mathbf{M}_{G}$ using TotalText [6] and TextSeg [51] datasets, both of which provide pixel-level annotations. Specifically, compared to conducting the cluster task on the whole image, we crop it into several text instances and then perform clustering to form $\mathbf{M}_{G}$. While the accuracy of the former is 32.35% for TotalText and 45.54% for TextSeg, $\mathbf{M}_{G}$ achieves 80.43% and 84.14%, respectively. Notably, the underlying morphological representations of glyphs exhibit relative invariance to slight structural changes, such as thicker or thinner. This reduces the dependence on pixel-level high-precision segmentation with expensive costs. Furthermore, although the clustering process introduces some noise, leading to inaccuracies in $\mathbf{M}_{G}$, it can still be viewed as a way of adding data noise in the image-mixing process, enhancing the robustness of text detection.